At 60°C, the reaction, PCl₅(g) ⇌ PCl₃(g) + Cl₂(g) has a K value of12.5. If the reaction quotient, Q, calculated from initial concentrations is 19.8 at 60°C, which of the following statements is correct?


there is more reactant concentration than product concentration


the rate of the reverse reaction is equal to the rate of forward reaction


the system is already at equilibrium


system must shift to the left to reach equilibrium


system must shift to the right to reach equilibrium


1
Expert's answer
2022-01-19T11:57:01-0500

Question:

At 60°C, the reaction, PCl₅(g) ⇌ PCl₃(g) + Cl₂(g) has a K value of12.5. If the reaction quotient, Q, calculated from initial concentrations is 19.8 at 60°C, which of the following statements is correct?


there is more reactant concentration than product concentration


the rate of the reverse reaction is equal to the rate of forward reaction


the system is already at equilibrium


system must shift to the left to reach equilibrium


system must shift to the right to reach equilibrium


Solution:

The reaction quotient "Q" and equilibrium constant "K" are calculated using similar expressions. For the equilibrium constant, equilibrium partial pressures are used:

"K_p = \frac{p(Cl_2)·p(PCl_3)}{p(PCl_5)} = 12.5".

For the reaction quotient, initial partial pressures are used instead:

"Q = \frac{p^0(Cl_2)·p^0(PCl_3)}{p^0(PCl_5)} = 19.8" .

From these expressions, one can see that when the reaction quotient is higher than the equilibrium constant, the concentration of the products must decrease in order to reach equilibrium. Therefore, system must shift to the left.


Answer: system must shift to the left to reach equilibrium